Mr.Akram invested Rr. 55000 in shares of two different companies A and B . company A announced a dividend 6% and company B of 8% .if the total income of the year received by Mr.Akram from these investment is Rs. 3900 .what were his investment in companies A and B ?
Answer by mananth(16946)

Part I 6.00% per annum ------------- Amount invested =x
Part II 8.00% per annum ------------ Amount invested = y
Total ---- 55000
Interest----- 3900

Part I 6.00% per annum ---x
Part II 8.00% per annum ---y
Total investment
x + 1 y= 55000 -------------1
Interest on both investments
6.00% x + 8.00% y= 3900
Multiply by 100
6 x + 8 y= 390000.00 --------2
Multiply (1) by -6
we get
-6 x -6 y= -330000.00
Add this to (2)
0 x 2 y= 60000
divide by 2
y = 30000
Part I 6.00% Rs 25000
Part II 8.00% Rs 30000